OpenStudy (anonymous):

-2(4n + 1) = 16 How do I solve for n?

OpenStudy (anonymous):

Divide by -2 on both sides

OpenStudy (anonymous):

4n + 1 = -8 4n = -9 n = -9/4

OpenStudy (anonymous):

i think that is that i maybe be wrong 5n -2=16 5n =14 5n/14 = n=0.35

OpenStudy (anonymous):

-2(4n + 1) = 16 -8n-2=16 -8n=18 n=-2.25 or 9/-4

OpenStudy (anonymous):

-9/4*